OpenStudy (anonymous):

Find the equation of the tangent line to curve y=x (squareroot) y at point (16,64)

OpenStudy (anonymous):

y = x^1/2 is eq of curve. dy/dx = slope of tangent = 1/2 x^(-1/2) We know that, eq of line is of the form y-y1/x-x1 = m Substitute values and ta-da! :)

OpenStudy (anonymous):

what exactly am i substituting into where sorry not very good at this aha! @ApoorvaAnand

OpenStudy (anonymous):

:) No problem dy/dx = 1/2 x^(-1/2) i.e. 1/2 * 1/4 (sub value of x) = 1/8 Now, we know x1 and y1 as 16 and 64. Therefore, eq of tangent = (x-16)/(y-64) = 1/8 Simplify to y = mx + c form if you want to.